AFN EQUATION

Carlsbad Corporation's sales are expected to increase from $5 million in 2016 to $6 million in 2017, or by 20%. Its assets totaled $2 million at the end of 2016. Carlsbad is at full capacity, so its assets must grow in proportion to projected sales. At the end of 2016, current liabilities are $1 million, consisting of $250,000 of accounts payable, $500,000 of notes payable, and $250,000 of accrued liabilities. Its profit margin is forecasted to be 4%, and the forecasted retention ratio is 45%. Use the AFN equation to forecast the additional funds Carlsbad will need for the coming year. Write out your answer completely. For example, 5 million should be entered as 5,000,000. Round your answer to the nearest cent.

$__________?
Answer #1

Given Expected sales= 6000000 After tax margin = 6000000 *4% = 240000 Dividend payout = 240000*(1-45%) 132000 Additions to re